OpenStudy (s):

Integral 2dx/ (x^2*(x^2+9)^(1/2)) ?

OpenStudy (anonymous):

Yesterday you were doing derivatives, now you're doing integrals?

OpenStudy (s):

yes, I tried to substitute x=3 tan theta, but then I got stuck later on..

OpenStudy (anonymous):

So you have: \[ \frac{2}{9\tan^2\theta \cdot 3\sec\theta} \]

OpenStudy (s):

there should be dx on top as well substituted with 2* 3 sec^2 theta d theta

OpenStudy (anonymous):

Wait so we have: \[ \frac{6\sec\theta}{27\tan^2\theta} \]

OpenStudy (anonymous):

Multiply top and bottom by \(\cos^2\theta\):\[ \frac{2\cos\theta}{9\sin^2\theta} \]

OpenStudy (anonymous):

\[ \cos\theta d\theta = d(\sin\theta) \]

OpenStudy (anonymous):

Meaning that \[ \frac{2\cos\theta\;d\theta}{9\sin^2\theta} = \frac{2\;d(\sin\theta)}{9\sin^2\theta}= \frac{2\;du}{9u^2} \]

OpenStudy (s):

I am a little confused how you got 6secθ/ 27tan2θ and then 2cosθ/ 9sin2θ

OpenStudy (s):

we got 2/ 9 tan ^2 theta * 3 sec theta.... did you expressed it in terms of sin and cos?

OpenStudy (anonymous):

Yes

OpenStudy (s):

I got 2 cos^3 theta/ 27 sin^2 theta

OpenStudy (anonymous):

\[ \int \frac{2dx}{x^2\sqrt{x^2+9}} \]

OpenStudy (anonymous):

\[ \begin{split} \int \frac{2\;dx}{x^2\sqrt{x^2+9}} &= \int \frac{2\;d(3\tan\theta)}{(3\tan\theta)^2\sqrt{(3\tan\theta)^2+9}}\\ &= \int \frac{6\sec^2\theta \;d\theta}{9\tan^2\theta|3\sec\theta|}\\ &= \int \frac{2\sec\theta \;d\theta}{9\tan^2\theta} \times\frac{\cos^2\theta}{\cos^2\theta}\\ &= \int \frac{2\cos\theta \;d\theta}{9\sin^2\theta} \\ &= \int \frac{2\;d(\sin\theta)}{9(\sin\theta)^2} \\ &= \int \frac{2\;du}{9u^2} \\ &= \frac{2}{9}\int u^{-2}\;du \\ &= \frac{2}{9}\int d\left(-u^{-1} \right)\\ &= -\frac{2}{9 u}+C \end{split} \]

OpenStudy (s):

So the final answer in terms of x would be -2(9+x^2)^(1/2) / 9x ??

OpenStudy (anonymous):

Well, \[ u = \sin\theta\\ x=3\tan\theta \]

OpenStudy (anonymous):

\[ \theta=\tan^{-1}\left(\frac{x}{3}\right) \]

OpenStudy (anonymous):

\[ u = \sin\left(\tan^{-1}\left(\frac x3\right)\right) \]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ u = \frac{x}{\sqrt{x^2+9}} \]

OpenStudy (anonymous):

So \[ -\frac {2}{9u}+C = -\frac{2\sqrt{x^2+9}}{9x}+C \]

OpenStudy (s):

that's what I got as well ! Thank you for double checking!